### v3

# A Geometric Algorithm for Tubular Shape Reconstruction from Skeletal Representation

Abstract
We introduce a novel approach for the reconstruction of tubular shapes from skeletal representations. Our method processes all skeletal points as a whole, eliminating the need for splitting input structure into multiple segments. We represent the tubular shape as a truncated signed distance function (TSDF) in a voxel hashing manner, in which the signed distance between a voxel center and the object is computed through a simple geometric algorithm. Our method does not involve any surface sampling scheme or solving large matrix equations, and therefore is a faster and more elegant solution for tubular shape reconstruction compared to other approaches. Extensive experiments demonstrate the efficiency and effectiveness of the proposed method. Code is avaliable at https://github.com/wlsdzyzl/Dragon.
Keywords Tubular structure  $\cdot$ Shape reconstruction  $\cdot$ TSDF  $\cdot$ Skeleton
1 Introduction
Shape reconstruction is a popular topic and has garnered significant attention from researchers due to its wide range of applications in computer graphics and computer vision. Existing methods focus on reconstructing object or scene from a point cloud or a sequence of image. Point cloud-based methods utilize a continuous high-dimensional function to fit the surface Carr et al. (2003); Kazhdan et al. (2006); Hou et al. (2022). These methods usually involve solving large linear equations and heavily rely on the quality of the point cloud and accurate estimation of surface normals. Image-based approaches iteratively fuse input frames into a distance or radiance field Nießner et al. (2013); Han and Fang (2018); Mildenhall et al. (2021); Dong et al. (2018). Accurate camera pose estimation is necessary to generate a consistent radiance field. The above methods usually require the use of cameras or radars to capture the surface of the object, and have the ability to perform high-quality reconstruction of the geometry and texture details. However, they struggle to generate complete and topologically correct models due to scan occlusion. Post-processing algorithms such as point cloud completion Huang et al. (2020); Wen et al. (2020); Kraevoy and Sheffer (2005); Dai et al. (2017) also have difficulty dealing with complex topological structures. As a consequence, these methods are not suitable for accurate shape reconstruction of tubular structures, particularly complex vascular trees, which are characterized by their intricate topology.
Skeletons serve as compact representations of primitive structures with identical topology Choi et al. (1997), and are inherently well-suited for effectively representing tubular structures. Skeletonization of tubular structures has undergone extensive research and demonstrated its utility across various applications Zwettler et al. (2008); Wang et al. (2020); Shit et al. (2021). However, recovering original shape from skeletal representation has been less studied. In practice, obtaining the skeleton or centerline of an object is much easier compared to acquiring the complete shape of the object. For example, annotating the centerline of coronary artery in CCTA requires much less labor compared to annotating the entire vessel Metz et al. (2008); Schaap et al. (2009). Shape reconstruction of centerlines can help generate binary mask volume and further improve the performance of segmentation model.
This study presents a novel approach for reconstructing tubular shapes from skeletal representations. In particular, each skeletal point is considered as a slice and the recovered shape is considered as a linear interpolation of these slices. Based on this simple assumption, we propose a pure geometric algorithm to compute the signed distance between any given location and the object. The space is represented as sparse voxels whose TSDF values are efficiently computed in parallel and stored in a spatial hashing table. Finally, we use a multi-threaded marching cubes algorithm to extract the isosurface as triangle meshes. In our approach, all skeletal points are processed simultaneously, thereby obviating the requirement for splitting the object into multiple segments. In addition, the proposed method does not involve any surface sampling scheme or solving large matrix equations, resulting in a faster and more elegant solution of tubular shape reconstruction.
2 Related Works
Tubular structure and skeleton representation: Analysis of tubular structure plays an important role in many fields of research ranging from civil engineering to medical imaging. Various works Xing and Makinouchi (2001); Hu et al. (2019); Luo et al. (2021); Gharleghi et al. (2022) are proposed for designing and segmenting tubular structures, for which the topology is their most important characteristic. Skeletons are concise representations of original structures with identical topology Choi et al. (1997) and are inherently well-adapted for efficiently depicting tubular structures. Sufficient research focus has been directed towards the skeleton extraction and application. Huang et al. (2013); Qin et al. (2019) extract curve skeleton from point clouds through medial axis transforms and mass-driven optimal transports. Lin et al. (2021) represents skeleton points as weighted summations of surface points and use a deep network to estimate the weights. Shit et al. (2021); Zhang et al. (2023a) introduce skeletonization in tubular structure segmentation and demonstrate its topology-preserving capability.
Implicit Representation for 3D Scene: Implicit scene representations, notably through Signed Distance Function (SDF), have become prominent in computer graphics and vision for efficiently capturing complex geometry. Lorensen and Cline (1987) firstly proposed a shape reconstruction algorithm from SDF by constructing triangles among 8 voxels. Kazhdan et al. (2006) utilizes Octree for accelerated solving of the Poisson equations to compute the indicator function. Recent studies have explored the implicit representation of continuous 3D shapes as level sets through deep networks. Park et al. (2019); Jiang et al. (2020) learns continuous SDF from point clouds. Mescheder et al. (2019); Genova et al. (2020) maps xyz coordinates to occupancy fields, which describes whether a point in 3D space is occupied by a surface or object. Mildenhall et al. (2021) represents scenes as neural radiance fields, which are widely used for novel view synthesis Barron et al. (2021); Pumarola et al. (2021), shape reconstruction Oechsle et al. (2021); Wang et al. (2021) and localization Rosinol et al. (2022); Deng et al. (2023).
Shape Reconstruction: Shape reconstruction tasks typically unfold in two distinct scenarios: reconstructing shapes from point clouds or reconstructing shapes from image sequences. For point cloud-based shape reconstruction, Bernardini et al. (1999) directly connects points to form triangles following a simple ball pivoting rule, without an continuous implicit representation. Carr et al. (2003) fits a radial basis function to represent the surface. Kazhdan et al. (2006); Hou et al. (2022) compute the indicator function by iteratively solving the Poisson equations. These methods try to compute an implicit representation, so that the surface can be further extracted through March Cube algorithms Lorensen and Cline (1987). A fine surface normal estimation is necessary for high reconstruction quality. Lin et al. (2022) views surface normals as unknown parameters in the Gauss formula to reconstruction shape without normals. Xu et al. (2023) propose a novel method for globally consistent normal estimation by regularizing the Winding-Number field. Image sequence-based approaches recover 3D shapes from 2D/3D frames instead of a whole point cloud, therefore accurate camera pose estimation is usually important. Nießner et al. (2013); Han and Fang (2018); Dong et al. (2018) perform camera pose estimation and incrementally fuse input frames into distance fields. Ji et al. (2017); Mildenhall et al. (2021) take camera parameters and images as input and generate a consistent implicit representation for scene rendering and reconstruction.
3 Method
Figure 1: Overview of our shape reconstruction algorithm. Left: original shape and sampled skeletal points. Middle: Constructed graph. Right: recovered shape from slices.
Our method tries to recover tubular shape from its skeletal representation that consists of a skeletal point set $\mathcal{C}=\{\mathbf{c}_{1},\mathbf{c}_{2},...,\mathbf{c}_{m}\}$ and the corresponding radius $\mathbf{r}\in\mathbb{R}^{m}$. Fig. 1 gives an overview of the reconstruction processing. Firstly, we need to know the connections among the input skeletal points, which is achieved through an adaptive graph construction algorithm as discussed in Sec. 3.1. Sec. 3.2 elaborates our geometric algorithm that enables fast computation of signed distance function (SDF). Finally, we introduce how to efficiently represent the shape through voxel hashing strategy and extract the dense model using march cubes algorithm in Sec. 3.3.
3.1 Graph Construction
Algorithm 1 Adaptive Graph Construction
skeletal points $\mathcal{C}$, number of KNN searched neighbors $k$
distance multiplier $m$ and angle threshold $a_{t}$
constructed graph $\mathcal{G}$
$\mathcal{G}\leftarrow$ new Graph $\triangleright$ Graph initialization
$\mathcal{G}.\mathcal{V}\leftarrow\mathcal{C}$
$T\leftarrow$ new KDTree($\mathcal{G}.\mathcal{V}$)
for $\mathbf{v}$ in $\mathcal{G}.\mathcal{V}$  do
     $\mathcal{I}\leftarrow T$.knn_search($\mathbf{v}$, $k$)
     $d_{\text{min}}\leftarrow$ dist$(\mathbf{v},\mathcal{G}.\mathcal{V}[\mathcal{I}[1]])$triangleright Calculate distance minimum, note that $\mathcal{I}[0]$ is the index of $\mathbf{v}$.
     $\mathcal{D}_{\text{edge}}\leftarrow\emptyset$triangleright Direction set of edges.
     for $i$ in $\mathcal{I}$ do
         $v_{i}\leftarrow\mathcal{G}.\mathcal{V}[i]$
         if dist$(\mathbf{v},\mathbf{v}_{i})>m\times d_{\text{min}}$ then
              break
         end if
         $f\leftarrow$ true $\triangleright$ Flag of new direction
         for $\mathbf{d}$ in $\mathcal{D}_{\text{edge}}$ do
              if $\text{angle}(\mathbf{v}_{i}-\mathbf{v},\mathbf{d})<=a_{t}$ then
                  $f\leftarrow$ false
                  break
              end if
         end $\triangleright$ If <$\mathbf{v},\mathbf{v}_{i}$> a new direction, add it into the edge set
         if $f$ is true then
              $\mathcal{G}.\mathcal{E}.\text{add}(<\mathbf{v},\mathbf{v}_{i}>)$
              $\mathcal{D}_{\text{edge}}.\text{add}(\mathbf{v}_{i}-\mathbf{v})$
         end if
     end for
end for
return $\mathcal{G}$
To reconstruct the shape, we need to obtain the connections among these skeletal points, which are inherently implied for the ordered skeletal points. However, ordered points are only suitable to represent a single curve without any loop or bifurcation. To recover the shape of a complex structure, we need to split the whole structure into multiple segments without loops or bifurcations and reconstruct all the segments separately, which is a tedious work. In this section, we propose an adaptive graph construction algorithm to find connections among unordered skeletal points based on their distance.
We construct an undirected graph $\mathcal{G}=<\mathcal{V},\mathcal{E}>$ whose vertices are the input skeletal points: $\mathcal{V}=\mathcal{C}$. Specifically, for each skeletal point $\mathbf{c}\in\mathcal{C}$ we find its $k$ nearest neighbors $\mathbf{c}_{1},\mathbf{c}_{2},...,\mathbf{c}_{k}$ that are sorted in ascending order based on their distances to point $\mathbf{c}$. We use $m$ and $a_{t}$ to denote the distance multiplier and angle threshold, respectively. The distance between $\mathbf{c}$ and $\mathbf{c}_{i}$ is denoted as $d_{\mathbf{c},\mathbf{c}_{i}}$. We establish an edge $<\mathbf{c},\mathbf{c}_{i}>$, under the condition that $d_{\mathbf{c},\mathbf{c}_{i}}$ does not exceed $m$ times the distance between $\mathbf{c}$ and it’s nearest neighbor $d_{\mathbf{c},\mathbf{c}_{1}}$, and the angles between $<\mathbf{c},\mathbf{c}_{i}>$ and the established edges are greater than $a_{t}$. This simple strategy obviates the necessity of selecting a suitable distance threshold for each input point cloud. It also avoids the constraint of k-nearest neighbors algorithm-based graph construction that each point need to be connected with a fixed number of points. Alg. 1 gives a detailed explanation of the proposed method.
Figure 2: Illustration of different skeletal representations. Left: original shape. Middle: the input skeletal points are a thinner representation of the original shape, but not a string-like structure. Right: the ideal skeletal points forms a string-like structure.
Radius-based clustering: In practice, the skeletal points we acquired may just exhibit a thinned representation of the original shape, and not necessarily be a string-like structure as shown in Fig. 2. This may lead to the construction of an extremely complex graph with a large number of redundant edges. To tackle this concern, we employ a pre-processing step that applies a radius-based clustering for the input skeletal points as summarized in Alg. 2. In specific, we randomly select a skeletal point $\mathbf{c}\in\mathcal{C}$ as a cluster center with corresponding radius $r$. The skeletal points that fall into the ball with center $\mathbf{c}$ and radius $s\cdot r$ are considered to come from the same cluster, which is implemented through radius searching based on KDTree Friedman et al. (1977) and $s$ is the searching strength. We continuously select cluster centers from the un-clustered points and repeat the process until all the skeletal points are successfully clustered. For each cluster, we compute the mean position and radius as a new skeletal point. Through these steps, we successfully reduce the time complexity without a significant decline in reconstruction quality.
Algorithm 2 Radius-based Clustering
skeletal points $\mathcal{C}$, radius $\mathbf{r}$ and strength $s$
clustered skeletal points $\mathcal{C}^{\prime}$ and radius $\mathbf{r}^{\prime}$
$\mathcal{C},\mathbf{r}\leftarrow\text{shuffle}(\mathcal{C},\mathbf{r})$.
$f\leftarrow\text{list}(|\mathcal{C}|,\textbf{false})$triangleright List of flags indicating whether elements are clustered
$T\leftarrow$ new KDTree($\mathcal{C}$)
$C^{\prime},r^{\prime}\leftarrow\emptyset$
[…]
     if $f[i]$ is true then
         continue
     end if
     $\mathcal{I}\leftarrow T$.radius_search($\mathcal{C}[i]$, $s\cdot\mathbf{r}[i]$) $\triangleright$ Indices of searched neighbors
     $\mathbf{c}\leftarrow\mathbf{0},r\leftarrow 0,n\leftarrow 0$
     for $j$ in $\mathcal{I}$ do
         if $f[j]$ is true then
              continue
         end if
         $f[j]\leftarrow$ true
         $\mathbf{c}\leftarrow\mathbf{c}+\mathcal{C}[j]$
         $r\leftarrow r+\mathbf{r}[j]$
         $n\leftarrow n+1$
     end for
     $\mathcal{C}^{\prime}$.add($\mathbf{c}/n$)
     $\mathbf{r}^{\prime}$.add($r/n$) $\triangleright$ New skeletal point is the center of a cluster
end for
return $\mathcal{C}^{\prime},\mathbf{r}^{\prime}$
3.2 SDF Computation
TSDF represents the geometry of a scene or object in a implicit way. For any point in the space, which is usually the center of a voxel, its signed distance function (SDF) value is defined as the signed distances to the nearest surface. If the point is located inside the object, the sign of distance is negative. TSDF is the SDF truncated by a distance threshold, so that we can only focus on voxels that are near to the surface. Because we use linear interpolation of slices to approximate the original shape, the SDF can be computed through a pure geometric way. Fig. 3 gives an illustration of SDF computation for a shape between two slices.
Figure 3: The geometric view of SDF computing.
Given two adjacent slices $S_{a}=(\mathbf{c}_{a},\mathbf{n}_{a},r_{a})$, $S_{b}=(\mathbf{c}_{b},\mathbf{n}_{b},r_{b})$ and a voxel center point $\mathbf{v}$, we use $f_{\text{SDF}}(S_{a},S_{b},\mathbf{v})$ to denote the SDF value of voxel $\mathbf{v}$ for the shape between slices $S_{a}$ and $S_{b}$, where $|a-b|=1$. The plane $P$ that is uniquely determined by $\mathbf{c}_{a}$, $\mathbf{c}_{b}$ and $\mathbf{v}$ intersects with $S_{a}$ and $S_{b}$ at four points, denoted as $\mathbf{i}_{1}$, $\mathbf{i}_{2}$, $\mathbf{i}_{3}$ and $\mathbf{i}_{4}$, respectively. The absolute value of $f_{\text{SDF}}(S_{a},S_{b},\mathbf{v})$ is the distance between $\mathbf{v}$ and the nearest surface, which is also the minimum distance between $\mathbf{v}$ and both line segments $\mathbf{i}_{1}\mathbf{i}_{4}$ and $\mathbf{i}_{2}\mathbf{i}_{3}$. The sign of $f_{\text{SDF}}(S_{a},S_{b},\mathbf{v})$ depends on whether the point $\mathbf{v}$ falls into the quadrilateral $\mathbf{i}_{1}\mathbf{i}_{2}\mathbf{i}_{3}\mathbf{i}_{4}$.
First, we compute the intersection points of slice $S_{a}$ and plane $P$. To simplify the computation, we transform the global to local coordinates where the origin is $\mathbf{c}_{1}$ and the direction vector describing z-axis is $\mathbf{n}_{a}$. We use $\mathbf{z}=(0,0,1)$ to denote the local coordinate of z-direction. The rotation matrix $\mathbf{R}$ from $\mathbf{z}$ to $\mathbf{n}_{a}$ can be computed by solving a linear equation:
$\mathbf{R}\cdot\mathbf{z}=\mathbf{n}_{a},\text{ subject to }\mathbf{R}\cdot\mathbf{R}^{\top}=\mathbf{I},$ (1)
where $\mathbf{I}$ is the identity matrix. Here we use a more efficient solution provided by Rodrigues (1840); Rasala (1981):
$\mathbf{R}=\mathbf{I}+\sin{\theta}\mathbf{K}+(1-\cos\theta)\mathbf{K}^{2}.$ (2)
In the above, $\mathbf{K}=[\mathbf{z}\times\mathbf{n}_{a}]_{\times}$ is a skew-symmetric matrix, $\theta$ is the rotation angle, while $\sin\theta=\|\mathbf{z}\times\mathbf{n}_{a}\|$ and $\cos\theta=\mathbf{z}\cdot\mathbf{n}_{a}$. The translation from origin $(0,0,0)$ to $\mathbf{c}_{a}$ is $\mathbf{c}_{a}$, therefore we can easily get the transformation matrix from local to global coordinates $\mathbf{T}=\left[\mathbf{R}|\mathbf{c}_{a}\right]$ whose inverse is the transformation matrix from global to local coordinates.
The normal vector $\mathbf{n}_{P}$ of plane $P$ can be computed through following equation:
$\mathbf{n}_{P}=\frac{(\mathbf{c}_{a}-\mathbf{v})\times(\mathbf{c}_{b}-\mathbf{v})}{\|(\mathbf{c}_{a}-\mathbf{v})\times(\mathbf{c}_{b}-\mathbf{v})\|}.$ (3)
The intersections $\mathbf{i}_{1}$ and $\mathbf{i}_{2}$ of plane $P$ and slice $S_{a}$ can be calculated by solving following system of equations:
$\mathbf{n}_{P}^{l}\cdot\mathbf{p}=0$ (4)
$\|\mathbf{p}\|=r_{a}$ (5)
$\mathbf{p}_{z}=0.$ (6)
Equ. 4 represents the plane $P$, where $\mathbf{n}_{P}^{l}=\mathbf{R}^{-1}\cdot\mathbf{n}_{P}$. Equ. 5 and Equ. 6 represent the slice $S_{a}$. The transformation from global to local coordinates actually degrades the solutions from to 3D into 2D, because all points have a z-axis value of zero. The solutions of above system are denoted as $\mathbf{i}_{1}^{l}$ and $\mathbf{i}_{2}^{l}$, respectively, and we have:
$\mathbf{i}_{1}^{l}=-\mathbf{i}_{2}^{l}.$ (7)
Transform the solutions back to global coordinates, we have:
$\mathbf{i}_{1}=\mathbf{T}\cdot\mathbf{i}_{1}^{l},\mathbf{i}_{2}=\mathbf{T}\cdot\mathbf{i}_{2}^{l}.$ (8)
The point that is on the same side of line $\mathbf{c}_{a}\mathbf{c}_{b}$ with $\mathbf{v}$ is denoted as $\mathbf{i}_{a}$, which is an end point of the line segment closer to $\mathbf{v}$.
Vice versa, we compute the $\mathbf{i}_{b}$ for slice $S_{b}$. The value of SDF can be calculated as follows:
$f_{\text{SDF}}(S_{a},S_{b},\mathbf{v})=\left\{\begin{matrix}-\text{dist}(\mathbf{i_{a}}\mathbf{i_{b}},\mathbf{v}),&\text{if }\mathbf{v}\text{ is in thequadrilateral }\mathbf{c}_{a}\mathbf{c}_{b}\mathbf{i}_{b}\mathbf{i}_{a};\\ \text{dist}(\mathbf{i_{a}}\mathbf{i_{b}},\mathbf{v}),&\text{otherwise}.\end{matrix}\right.$ (9)
where $\text{dist}(\cdot,\cdot)$ returns the distance between a line segment and and a point.
The function $f_{\text{SDF}}(S_{a},S_{b},\mathbf{v})$ calculates the SDF of voxel $\mathbf{v}$ for partial shape between two adjacent slices $S_{a}$ and $S_{b}$. Given graph $\mathcal{G}$, the SDF of $\mathbf{v}$ for the whole structure is defined as:
$f_{\text{SDF}}(\mathbf{v})=\min\{\min_{e\in\mathcal{G.E}}f_{\text{SDF}}(S_{e[0]},S_{e[1]},\mathbf{v}),d_{\text{ball}}\},$ (10)
where $e[0]$ and $e[1]$ is the indices of two endpoints of edge $e$, $d_{\text{ball}}$ is a smoothing factor, which is defined as the signed distance to the ball determined by the nearest skeletal point $\mathbf{c}_{n}$: $d_{\text{ball}}=\|\mathbf{v}-\mathbf{c}_{n}\|-r_{n}$.
Figure 4: A failure case when faced with poor sampling.
To obtain the SDF of $\mathbf{v}$ for the whole shape, we need to compute SDF values of $\mathbf{v}$ for all edges in the graph, which is time-consuming and inefficient. In our implementation, we only compute the SDF values for 5 nearest edges. This is based on a simple assumption: the skeletal points is well sampled so that far skeletal points have little effect on the SDF computing. Fig. 4 shows a failure case if the assumption is not satisfied, in which all 5 nearest skeletal points of a voxel inside the tube are outside the corresponding partial shape, leading to an incorrect SDF computation.
Fast SDF Approximation: The aforementioned method computes the distance between a voxel center and it’s nearest surface that is denoted as $d$ in Fig. 3. A more fast way to approximate the SDF value for a shape between two adjacent slices $(\mathbf{c}_{a},\mathbf{n}_{a},r_{a})$ and $(\mathbf{c}_{b},\mathbf{n}_{b},r_{b})$ is to assume $\mathbf{n}_{a}=\mathbf{n}_{b}=\mathbf{c}_{b}-\mathbf{c}_{a}$ and compute $d_{\text{f}}$. For a given voxel center $v$, we project it to line segment $\mathbf{c}_{a}\mathbf{c}_{b}$ at a new skeletal point $\mathbf{c}_{v}$, whose radius can be calculated by a simple interpolation:
$r_{v}=\frac{r_{b}\cdot\|\mathbf{c}_{v}-\mathbf{c}_{a}\|+r_{a}\cdot\|\mathbf{c}_{b}-\mathbf{c}_{v}\|}{\|\mathbf{c}_{b}-\mathbf{c}_{a}\|}.$
And then we have:
$d_{\text{f}}=\|\mathbf{v}-\mathbf{c}_{v}\|-r_{v}.$
Figure 5: Surface extraction using different SDF values.
Note that the approximated SDF value calculated through fast computation still generates an accurate surface under the equal normal assumption. As shown in Fig. 5, $d_{1}$ and $d_{2}$ represent the SDF values of two adjacent voxel centers $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$, respectively, while $d_{f1}$ and $d_{f2}$ represent the approximated SDF values. According to the principle of similar triangles, we have $\frac{d_{1}}{d_{2}}=\frac{d_{f1}}{d_{f2}}$. Because the surface point is the weighted interpolation of voxel centers which can be formulated as:
$\mathbf{v}=\frac{d_{2}}{d_{1}+d_{2}}\cdot\mathbf{v}_{1}+\frac{d_{1}}{d_{1}+d_{2}}\cdot\mathbf{v}_{2}=\frac{1}{\frac{d_{1}}{d_{2}}+1}\cdot\mathbf{v}_{1}+\frac{1}{1+\frac{d_{2}}{d_{1}}}\cdot\mathbf{v}_{2},$
using the approximated SDF values generates the same result as using true SDF values. Note that, providing an algorithm to compute real SDF values is not without merit, as real SDF values are more useful in some tasks Park et al. (2019).
3.3 Voxel Hashing and Mesh Extraction
Instead of computing the SDF value for each voxel in the bounding box, we use a spatial hashing table Nießner et al. (2013); Han and Fang (2018) to store sparse cubes that near to the object, which greatly reduces the required memory and computational complexity. The space is represented as cubes, while each cube consists of $8\times 8\times 8$ voxels. We compute a bounding box for the input skeletal points, any cube that is located within the bounding box or has an overlapping region with the bounding box is referred to a candidate cube. For each candidate cube, we calculate the SDF values for its 8 corner voxels. If the minimum absolute value among these values is smaller than the truncated distance, or if the signs of these values are not the same, we consider that there might be some parts of surface in the cube, and the cube is added into the hash table. The truncated SDF is then calculated for all voxels in selected cubes. Finally, we use the multi-threaded marching cubes algorithm Lorensen and Cline (1987) to extract triangle meshes. The corresponding mask volume can also be easily obtained from the TSDF by identifying whether the voxels are located inside the object. Overall, the proposed shape reconstruction algorithm is summarized in Arg. 3.3. The default values of mentioned parameters during the processing are listed in Tab. 1.
Algorithm 3 Shape reconstruction from unordered skeletal points
skeletal points $\mathcal{C}$, radius $\mathbf{r}$, number of KNN searched neighbors $k$
clustering strength $s$, voxel resolution $l$ and truncated distance $d_{t}$
distance multiplier $m$, angle threshold $a_{t}$
$M$: reconstructed triangle mesh
$\mathcal{C},\mathbf{r}\leftarrow\text{radius\_clustering}(\mathcal{C},\mathbf{r},s)$
$\mathcal{G}\leftarrow\text{adaptive\_graph\_construction}(\mathcal{C},k,m,a_{t})$
function (compute_sdf)$(\mathbf{v})$: $\triangleright$ SDF computation
     $\mathcal{I}\leftarrow T$.knn_search($\mathbf{v}$, $k$)
     $f_{\text{min}}\leftarrow d_{t}$
     for $i$ in $\mathcal{I}$ do
         $\mathbf{v}_{i}\leftarrow\mathcal{G}.\mathcal{V}[i]$
         for $\mathbf{v}_{j}$ in $\mathcal{G}.\mathcal{E}(\mathbf{v}_{i})$ do
              $f\leftarrow\text{compute\_sdf}(\mathbf{v},\mathbf{v}_{i},\mathbf{v}_{j})$triangleright As mentioned in Sec. 3.2
              if $f<f_{\text{min}}$ then
                  $f_{\text{min}}\leftarrow f$
              end if
         end for
     end for
     $d_{c}\leftarrow\text{dist}\left(\mathbf{v},\mathcal{G}.\mathcal{V}\left[\mathcal{I}[0]\right]\right)-\mathbf{r}\left[\mathcal{I}[0]\right]$triangleright Distance between $\mathbf{v}$ and the surface of the closest ball
     if $f_{\text{min}}>d_{c}$ then
         $f_{\text{min}}\leftarrow d_{c}$
     end if
     return $f_{\text{min}}$
end function
[…]
for $v$ in $\mathcal{H}$ do
     $H[\mathbf{v}]$.sdf $\leftarrow$ compute_sdf($\mathbf{v}$)
end for
$M\leftarrow$$\mathcal{H}$) $\triangleright$ Mesh extraction through Marching Cubes algorithm
return $M$
Table 1: Parameter table. Default values are set empirically.
Algorithm Parameter Notation Default Value
Radius clustering Clustering strength $s$ 0.75
Graph reconstruction Number of KNN searched neighbors $k$ 5
Distance multiplier $m$ 2.5
Angle threshold $a_{t}$ 75
SDF computation Number of KNN searched neighbors $k$ 5
Voxel resolution $l$ 0.025
Truncated distance $d_{t}$ 0.1
4 Experimental Results
4.1 Datasets
Figure 6: Qualitative experimental results. From top to bottom each row shows reconstruction results for datasets CAT08, ImageCAS, ATM22, PARSE2022 and DeepVesselNet, respectively. Columns from left to right are visualizations of skeletal points and reconstructed shapes using different methods. The recovered shapes by our method are more complete and accurate, achieving the best reconstruction quality.
We evaluate our method on five public datasets. CAT08 Schaap et al. (2009) is a coronary artery center-line extraction dataset which contains 8 CCTA volumes with carefully labeled center-lines. ImagesCAS Zeng et al. (2022) is a large coronary artery segmentation dataset that contains 1000 3D images with sizes of 512 $\times$ 512 $\times$ (206-275). ATM22 Zhang et al. (2023b) is a pulmonary airway segmentation dataset and contains 500 CT scans with sizes of 512 $\times$ 512 $\times$ (84-879). PARSE2022 Maurya et al. (2022) contains 200 3D volumes with sizes of 512 $\times$ 512 $\times$ (228-376) for pulmonary artery segmentation. DeepVesselNet Tetteh et al. (2020) contains 100 synthetic 3D images for complicated vessel segmentation. The images have a size of 325 $\times$ 304 $\times$ 600 voxels. All these segmentation datasets only contains labeled masks, and their skeletons are calculated through a binary thinning algorithm Zhang and Suen (1984).
Table 2: Quantitative experimental results.
Dataset Method Dice Score Radius Difference Center Agreement Average Running Time (s)
CAT08 (1496) Ball Pivoting - 0.2286 0.9194 0.18
Poisson - 0.1887 0.9999 0.61
Ours - 0.1013 1.0000 0.54
Ours (fast) - 0.1010 1.0000 0.24
ImageCAS (1507) Ball Pivoting 0.7007 0.7619 0.6999 0.22
Poisson 0.7548 0.4614 0.9939 0.76
Ours 0.7566 0.5983 0.9994 0.39
Ours (fast) 0.7638 0.5511 0.9997 0.22
ATM22 (2688) Ball Pivoting 0.3619 1.2396 0.7040 2.11
Poisson 0.6805 0.7867 0.9929 1.89
Ours 0.6882 0.4884 0.9992 1.56
Ours (fast) 0.7045 0.4048 0.9997 0.93
PARSE2022 (8885) Ball Pivoting 0.2810 0.8912 0.7527 8.50
Poisson 0.6263 0.6415 0.9915 4.87
Ours 0.6347 0.4883 0.9985 4.63
Ours (fast) 0.6516 0.3773 0.9994 3.21
DeepVesselNet (62519) Ball Pivoting 0.2960 0.6602 0.7943 73.32
Poisson 0.7387 0.3239 0.9987 32.13
Ours 0.7473 0.3394 0.9998 25.96
Ours (fast) 0.7498 0.2819 0.9999 19.54
4.2 Evaluation Metrics
We use three metrics to measure the reconstruction accuracy. We recover mask volume from reconstructed shape and compute Dice score between ground-truth and recovered mask for segmentation datasets. Radius difference is defined as the average different between the re-computed radius of recovered shape and original radius. Center agreement calculates the percentage of skeletal points that are located inside the recovered shape. Additionally, we compare the efficiency of different reconstruction algorithms through the average running time.
4.3 Qualitative and Quantitative Analysis
We compare our algorithm with two previous works. Ball Pivoting Bernardini et al. (1999) tries to connect the surface points to form triangles, while Screened Poisson Kazhdan and Hoppe (2013) calculates the indicator function by solving the regularized Poisson equation. Both of them are shape reconstruction algorithms from surface points and cannot be directly applied on skeletal representations. Therefore, for each slice we generate 10 surface points to run the baseline algorithms. Note that the slices are generated from constructed graph as mentioned in Sec. 3.1.
Tab. 2 shows the quantitative results, where we introduce the accuracy and performance of our reconstruction algorithms using real and approximated sdf values, denoted as ours and ours (fast), respectively . Our methods demonstrate high superiority in reconstruction accuracy. While the mask and shape recovered by ours (fast) maintains the highest consistency with original mask and skeleton, the reconstruction accuracy with real sdf values slightly decreases, possibly due to the precision loss incurred during the computation process. Meanwhile, the proposed methods are generally more efficient. Ours (fast) is 2.38$\times$ faster than Ball Pivoting Bernardini et al. (1999) and 1.67$\times$ faster than Screened Poisson Kazhdan and Hoppe (2013), respectively. It’s worth noting that the shapes recovered by our methods also have much smaller volumes than the ones by Screened Poisson Kazhdan and Hoppe (2013) while maintaining a higher reconstruction quality.
5 Conclusion
We propose a novel algorithmic pipeline to reconstruct tubular shape from skeletal representation. Specifically, we use a radius-based clustering as pre-processing to reduce point redundancy and obtain string-like skeletal points. Based on the simple assumption that tubular shape can be recovered through a linear interpolation of slices, we analyze the geometric solution of SDF computing and propose a fast SDF approximation. Finally, we use a paralleled marching cube algorithm to extract surface models. Experiments shows that our method achieves superior reconstruction quality with lower computational complexity compared to other methods.
Acknowledgements: This work is supported in part by the Natural Science Foundation of China (Grant 62371270) and Shenzhen Key Laboratory of Ubiquitous Data Enabling (No.ZDSYS20220527171406015).